# Persona-DB: Efficient Large Language Model Personalization for Response Prediction with Collaborative Data Refinement

### v1

@@ 3.3 General Case Results Discussion
We evaluate the effectiveness of the Persona-DB framework and the baselines, setting the top-k as 40 for both our variants and the baselines in the retrieval-augmentation step. From the table, it is evident that H-Recency and H-Retrieval underperform compared to History (Full), as expected, due to their less comprehensive view of the user’s information. On the other hand, Persona-DB consistently outperforms these retrieval-based methods across all evaluation metrics with a large margin in correlation, demonstrating that our framework can achieve superior performance under the same retrieval capacity, compared to using historical posts only. The table also indicates that while leveraging a user’s complete history offers a broad view of user data during inference, our approach using only the top-k personas, in fact, exceeds the performance of History (Full) consistently. This suggests that our strategy of self-improving and composing interrelated personas achieves better personalization efficiency as it uses much less data while achieving better results.
To underscore the significance of the collaborative stage within our framework further, we examine their absence’s effect on performance. Omitting JOIN operations results in a marked decrease in correlation metrics and F1 scores, affirming its essential role in addressing knowledge gaps among users. Interestingly, even without implementing JOIN, the hierarchical database exhibits greater predictive capability than that of the baselines using a log-only database. This observation reiterates the value of high-level insights in improving generalizability in the downstream models. While the hierarchical construction is not perfect due to the use of LLM, it still allows both components of our framework to work. This indicates that the method is robust against error propagation at some level, and further enhancements to the LLM to generate higher-quality insights may yield improved performance for Persona-DB.
In short, the results reveal that the Persona-DB framework, with its hierarchical persona database and collaborative refinement process, offers a more efficient and accurate solution for personalized LLM interactions. The critical role of the JOIN operation also shows the importance of developing methods that consider user relations for LLM personalization.
@@
Case Study: Lurkers
@@
Case Study: Frequent Users
@@
3.4 Results on Lurkers and Frequent Users
In our evaluation, we delve into the evaluation of Persona-DB across two types of common yet challenging cases: Lurkers and Frequent Users. This segmentation allows us to understand the framework’s robustness and performance in scenarios representing the extremes of user interaction volumes.
Lurkers   The term "Lurker" refers to users with minimal interaction history, a common scenario in recommendation systems known as the cold-start problem. Being capable of personalizing responses for lurkers indicates a framework’s ability to offer tailored interactions to new users without the prerequisite of extensive historical accumulation. We select 100 users with the sparsest (nonempty) interaction records, averaging only 13.81 interactions per user, for this study. The results, as illustrated in Table 2, highlight Persona-DB’s superior performance across all metrics compared to the baselines. Specifically, our framework shows a 15% improvement in Pearson correlation over the baseline method, showcasing its exceptional ability to maintain accuracy with minimal user data. This efficiency can be attributed to the framework’s capability to discern user characteristics from limited interactions, leveraging reliable, frequent collaborators’ insights to enhance personalization accuracy. Remarkably, even before applying the JOIN operation, the transformed database exhibits noticeably improved performance, underscoring the predictive power of LLMs in identifying generalizable personas.
@@ 3.4 Results on Lurkers and Frequent Users
The conducted analyses on both lurkers and frequent users highlight the versatility and robustness of the Persona-DB framework in delivering personalized LLM interactions. By handling the challenges posed by both sparse and extensive user interaction histories, our approach demonstrates a clear improvement in LLM personalization over the baselines. The effectiveness of collaborative persona databases across different scenarios shows the potential for practical applicability in real-world contexts.
Figure 3: Both variants of our framework are capable of maintaining performance above that of full history, even when the retrieval size is significantly reduced.
3.5 Personalization Efficiency Analysis
To minimize resource consumption in the inference stage, it is crucial to improve data efficiency in retrieval (i.e., the retrieved set remains generalizable while being reduced in size). This section delves deeper into this perspective by illustrating how our framework’s performance varies with different retrieval sizes.
Figure 3 demonstrates the robustness of our method, illustrating that accuracy is maintained even as the retrieval cap is significantly reduced. This underscores the method’s resilience to variations in context window size for LLMs, highlighting its capability to deliver precise predictions with minimal data retrieval. The compelling performance, even with a constrained database, supports our hypothesis that the personas derived from our hierarchical construction possess enhanced predictive power.
While larger retrieval capacity generally corresponds to improved performance, likely owing to the richer information context, certain instances reveal a slight decrease in metrics. This phenomenon aligns with existing researchs Wang et al. (2023) suggesting that overly extensive contexts can, at times, introduce more noise than they mitigate. This finding is intriguing as it shows there may exist an optimal retrieval threshold depending on the query, beyond which the additional data may become counterproductive. Moreover, we can see that as the retrieval size decreases, the advantage of using collaborative entries shrinks. This trend likely stems from the increased significance of a user’s own data in situations where retrieval capacity is limited.
@@ 4 Related Work
Personalization plays a crucial role in the effectiveness of large language models (LLMs), particularly in ensuring their alignment with individual preferences and interests Hwang et al. (2023); Kirk et al. (2023); Mao et al. (2023); Han et al. (2023). With the advent of LLMs, the landscape of personalization has also been significantly transformed, enabling enhanced user engagement and the delivery of services tailored to individual needs Chen et al. (2023b) and cultural preferences Li et al. (2023); Fung et al. (2023, 2024). Traditional personalization approaches for language models often use auxiliary encoders to learn user representation from user profile data King and Cook (2020); Qiu et al. (2021); Lian et al. (2022). For instance, to personalize pre-trained language models, UserIdenitifer Mireshghallah et al. (2022) prepends a user identifier prefix at the beginning of the input text during the fine-tuning process. Works have also been done to utilize tensor factorization during training to tackle data sparsity problems Wang et al. (2021). Yet such works would require training resources and can be challenging in situations of extensive user data and large models. Recently, there have been works focusing on utilizing and improving retrieval augmentation due to its capability to efficiently borrow external knowledge lacking in the pre-trained data Wang and McAllester (2020); Lewis et al. (2020) without the cost of fine-tuning downstream LLMs. In Zhang et al. (2023) and Asai et al. (2022), the authors introduced new embedding models to better support the diverse retrieval augmentation needs. Furthermore, Srinivasan et al. (2022) introduced a practical approach for distilling retrieval-augmented LLMs, and Guu et al. (2020) demonstrated the effectiveness of retrieval-augmented language model pre-training for open-domain question answering. Yet limited exploration has been done on optimizing the data to be retrieved to improve personalization. In our work, we are the first to introduce a collaborative-style strategy to enhance data efficiency in personalized retrieval for LLMs.
@@ 5 Conclusions and Future Work
In conclusion, we introduce Persona-DB, a framework designed to enhance the efficiency and accuracy of retrieval-based LLM personalization through collaborative data refinement. The framework leverages a hierarchical approach to user data representation, enabling the construction of insights that are generalizable across task contexts, and introduces a collaborative mechanism to effectively bridge communal knowledge for improved query response when the user lacks sufficient relevant data. Our experimental results indicate a marked improvement in performance, particularly in scenarios characterized by sparse user data, thus addressing a critical challenge in LLM personalization. Furthermore, we perform analysis to study the impact of data composition on retrieval effectiveness. Future work will focus on developing methods that refine the framework’s capability to dynamically adapt the matching process based on user feedback and interaction patterns.
@@
Our user retrieval process currently fetches cached personas from the highest and the most concise level of the hierarchy. However, conciseness implies a trade-off between importance and information richness. Hence, the process can be enhanced by selecting personas from different granularity levels depending on the task domains. For instance, we can leverage both users’ values and social roles, and slightly lengthier but more informative induced personas such as patterns analysis and opinions. Furthermore, we employ LLMs for abstracting personas in our work. However, just as data annotators may make errors, LLMs inevitably may overlook critical information or hallucinate incorrect viewpoints. We believe that this shortfall can be addressed through continuous improvement of the LLM’s capabilities.
Ethics
There are no ethics-related concerns in our paper. The dataset and resources come from publicly available repositories. Our framework is designed to enhance personalization efficiency, assisting humans in finding the most suitable content and information. Moreover, the effectiveness of our method on the news response prediction task shows its potential for improving personalized content moderation and contributing to the creation of a safer online environment.
@@ Acknowledgement
This research is based upon work supported in part by U.S. DARPA INCAS Program No. HR001121C0165. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of DARPA, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein. We would also like to thank Kevin Small for his helpful suggestions.
@@ A.1 Additional Details
The data statistics are shown in Table 3. In the main and case studies, we set 25% as the composition ratio. We show all prompts used in the work in the following figures. Figure 5 and Figure 6 represent the prompts for constructing the hierarchical database. Figure 7, Figure 8, and Figure 9 represent the inference prompt for baseline, Persona-DB w/o JOIN, and Persona-DB.
Figure 7: Prompt template used for Baseline Inference.
Figure 9: Prompt template used for Inference for Persona-DB.
@@
Table 3: Summary statistics for the original dataset.

### v3

@@ 3.3 General Case Results Discussion
We evaluate the effectiveness of the Persona-DB framework and the baselines, setting the top-k as 40 for both our variants and the baselines in the retrieval-augmentation step. From the evaluation on RFPN (Table 1), it is evident that H-Recency and H-Retrieval underperform compared to History (Full), as expected, due to their less comprehensive view of the user’s information. On the other hand, Persona-DB consistently outperforms these retrieval-based methods across all evaluation metrics with a large margin in correlation, demonstrating that our framework can achieve superior performance under the same retrieval capacity, compared to using historical posts only. The table also indicates that while leveraging a user’s complete history offers a broad view of user data during inference, our approach using only the top-k personas, in fact, exceeds the performance of History. This suggests that our strategy of self-improving and composing interrelated personas achieves better personalization efficiency as it uses much less data while achieving better results.
To underscore the significance of the collaborative stage within our framework further, we examine their absence’s effect on performance. Omitting JOIN operations results in a marked decrease in correlation metrics and F1 scores, affirming its essential role in addressing knowledge gaps among users. Interestingly, even without implementing JOIN, the hierarchical database exhibits greater predictive capability than that of the baselines using a log-only database. This observation reiterates the value of high-level insights in improving generalizability in the downstream models. We also observe that removing the intermediate persona layers (i.e. IP or DP), leads to a decrease in performance, underscoring their importance in enhancing the model’s capability. Additionally, we conduct a validation study to assess the quality of the LLM analyzer in stage 1 (detailed in the Appendix). Although the extraction is not perfect due to the use of LLM, it still allows both components of our framework to work. This indicates that the method is robust against error propagation at some level, and further enhancements to the LLM to generate higher-quality insights may yield improved performance for Persona-DB.
We also present results on OpinionQA in Figure 3. The plot reveals a similar pattern where Persona-DB surpasses the baseline models across topics. Notably, we see that compared with the baselines, it performs particularly well in the Global Altitudes category, consisting of questions on international affairs. This enhanced performance can be attributed to the importance of opinions from users within similar demographic groups in predicting current users’ opinions on this topic.
Figure 3: Performance comparison on the OpinionQA task. The plot shows that Persona-DB outperforms the baselines consistently.
@@
Lurkers (Cold-Start)
@@
w/o JOIN 47.07 53.34 73.79 49.79
Frequent Users
@@
3.4 Analysis on Lurkers and Frequent Users
In our evaluation, we delve into the evaluation of Persona-DB across two types of common yet challenging cases: Lurkers and Frequent Users. This segmentation allows us to understand the framework’s robustness and performance in scenarios representing the extremes of user interaction volumes. We perform the analysis on the RFPN task as it features high diversity in the history lengths.
Lurkers   The term "Lurker" refers to users with minimal interaction history, a common scenario in recommendation systems known as the cold-start problem. Being capable of personalizing responses for lurkers indicates a framework’s ability to offer tailored interactions to new users without the prerequisite of extensive historical accumulation. We select 100 users with the sparsest (nonempty) interaction records, averaging only 13.81 interactions per user. The results, as illustrated in Table 2, highlight Persona-DB’s superior performance across all metrics compared to the baselines. Specifically, our framework shows a 11% improvement in Pearson correlation over the baseline method, showcasing its exceptional ability to maintain accuracy with minimal user data. This efficiency can be attributed to the framework’s capability to discern user characteristics from limited interactions, leveraging reliable, frequent collaborators’ insights to enhance personalization accuracy. Remarkably, even before applying the JOIN operation, the transformed database exhibits noticeably improved performance, underscoring the predictive power of LLMs in identifying generalizable personas.
@@ 3.4 Analysis on Lurkers and Frequent Users
Top-K Method Pearson Accuracy
40 Persona-DB 47.88 62.66
40 IntSum 45.05 59.96
40 Hist-Recency 41.49 58.71
30 Persona-DB 47.45 61.89
30 IntSum 41.93 59.38
30 Hist-Recency 42.08 59.58
10 Persona-DB 45.46 60.35
10 IntSum 45.33 60.15
10 Hist-Recency 42.21 59.19
Table 3: Our framework is capable of maintaining performance above that of baselines, even when the retrieval size is significantly reduced.
3.5 Analysis on Varying Retrieval Sizes
To minimize context window cost in the inference stage, it is crucial to improve data efficiency in retrieval (i.e., the retrieved set remains generalizable while being reduced in size). This section explores how varying retrieval sizes affect our framework’s performance.
Table 3 demonstrates the robustness of our method, illustrating that accuracy is maintained above the baselines even as the retrieval cap is significantly reduced. This underscores the method’s resilience to variations in context window size for LLMs, highlighting its capability to deliver precise predictions with minimal data retrieval. The compelling performance, even with a constrained database, supports our hypothesis that the personas derived from our hierarchical construction possess enhanced predictive power. Moreover, we can see that as the retrieval size decreases, the advantage of using collaborative entries shrinks. This trend likely stems from the increased significance of a user’s own data in situations where retrieval capacity is limited.
@@ 4 Related Work
Personalization plays a crucial role in the effectiveness of large language models (LLMs), particularly in ensuring their alignment with individual preferences and interests Hwang et al. (2023); Kirk et al. (2023); Mao et al. (2023); Han et al. (2023); Chan et al. (2021); Ge et al. (2024); Salemi et al. (2023). With the advent of LLMs, the landscape of personalization has also been significantly transformed, enabling enhanced user engagement and the delivery of services tailored to individual needs Chen et al. (2023b) and cultural preferences Li et al. (2023b); Fung et al. (2023, 2024); Zhang et al. (2024). Traditional personalization approaches for language models often use auxiliary encoders to learn user representation from user profile data King and Cook (2020); Qiu et al. (2021); Wang et al. (2021); Li et al. (2023a). For instance, to personalize pre-trained language models, UserIdenitifer Mireshghallah et al. (2022) prepends a user identifier prefix at the beginning of the input text during the fine-tuning process. Yet such works would require training resources and can be challenging in situations of extensive user data and large models. Recently, there have been works focusing on utilizing and improving retrieval augmentation due to its capability to efficiently borrow external knowledge lacking in the pre-trained data Wang and McAllester (2020); Lewis et al. (2020) without the cost of fine-tuning downstream LLMs. In Zhang et al. (2023) and Asai et al. (2022), the authors introduced new embedding models to better support the diverse retrieval augmentation needs. Furthermore, Srinivasan et al. (2022) introduced a practical approach for distilling retrieval-augmented LLMs, and Guu et al. (2020) demonstrated the effectiveness of retrieval-augmented language model pre-training for open-domain question answering. Yet limited exploration has been done on optimizing the data to be retrieved to improve personalization.
@@ 5 Conclusions and Future Work
In conclusion, we introduce Persona-DB, a framework designed to enhance the accuracy and context efficiency of retrieval-based LLM personalization through collaborative data refinement. The framework leverages a hierarchical approach to user data representation, enabling the construction of insights that are generalizable across task contexts, and introduces a collaborative mechanism to effectively bridge communal knowledge for improved query response when the user lacks sufficient relevant data. Our experimental results indicate a marked performance improvement, particularly in scenarios characterized by sparse user data, thus addressing a critical challenge in LLM personalization. Furthermore, we perform analysis to study the impact of data composition on retrieval effectiveness. Future work will focus on developing methods that refine the framework’s capability to dynamically adapt the matching process based on user feedback and interaction patterns.
@@
We employ LLMs for abstracting personas in our work. Due to inherent bias and imperfection in the LLM, the quality of the extraction in hierarchical refinement can also affect the generalizability on diverse personas. We believe that this shortfall can be addressed through continuous improvement of the LLM’s capabilities. Moreover, in this study, our focus primarily lies on improving retrieval efficiency in downstream task scenarios, where we evaluate the model’s capability to preserve accuracy while reducing retrieval capacity, which in turn reduces the cost associated with online inference. In the hierarchical refinement stage, there is an existence of a one-time inference cost associated with preprocessing user data, a trade-off for enhancing data efficiency in downstream usage. We believe that using a smaller distilled model will alleviate the cost for this stage.
Ethics Statements
Our personalization approach involves storing and retrieving information about user preferences and personalities, which requires careful safety protocol and ethical consideration when deployed. While the extraction of generalized user personas and collaborative data composition offers powerful personalization capabilities, the sharing and combining of information between users requires consideration of privacy implications. The process should be implemented with appropriate privacy protections such as data anonymization, secure storage, and user consent.
Additionally, we recognize important considerations for responsible development: First, personas can risk oversimplifying and stereotyping users (as they are caricatures of people), potentially leading to biased or unrealistic representations that are unable to capture the true complexity and diversity of human behavior. Second, personalization technology enables detailed user profiling that could be misused for surveillance or targeting of individuals and groups. To mitigate these risks, the developers need to emphasize privacy-preserving techniques and aim to maintain user agency by being transparent about data usage.
Moreover, the datasets we used in this work are from publicly accessible repositories from existing publications. We only provide pointers to these existing data repositories and the data loading script that contains no user information. We will not share the datasets themselves. We will also not release any intermediate representations generated by models which may contain user information.
@@ Acknowledgement
This research is based upon work supported in part by U.S. DARPA INCAS Program No. HR001121C0165. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of DARPA, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.
@@ A.1 Additional Details
Our experiments were conducted using the PyTorch framework Paszke et al. (2019) and the Huggingface Transformers library Wolf et al. (2020). The sentiment intensity labels adhere to the definition in SemEval-2018 Task 1 Mohammad et al. (2018), incorporating both magnitude and direction in the evaluation of sentiment intensity. In our work, we use concatenation of histories to implement the JOIN in Equation 3 after collaborative databases have been retrieved. Future work will explore condensing the collaborative databases using LLMs.
The data statistics for RFPN are shown in Table 4. For OpinionQA, each topic we used contains a test split of 1000 samples. In particular, Biomedical and Food Issues contains 67 questions in the survey and there are 2537 respondents in total, Global Attitudes contains 104 questions and 2596 respondents, and America in 2050 contains 90 questions with 2524 respondents in total.
We show all prompts used in the work in the following figures. Figure 7 and Figure 8 represent the prompts for constructing the hierarchical database. Figure 9, Figure 10, and Figure 11 represent the inference prompt for baseline, Persona-DB w/o JOIN, and Persona-DB.
A.2 Validation Study on LLM Extraction
As the extracted personas in stage 1 serve as an intermediate layer to the method, and the LLM analyzer itself can introduce errors, we conducted a validation study regarding the quality of the extraction. Specifically, we sampled 50 LLM extraction results from user histories and distributed them to three human raters (who are graduate students who passed an initial quiz of 8 samples) to verify the accuracy. On average, the raters assigned a score of 3.9/5, indicating a notable level of precision despite imperfections in the LLM’s output. While the extraction accuracy isn’t imperfect, our results show that the quality of the current annotation does have the merit of contributing to performance positively.
A.3 Comparison with Prompt Compression
In our method, we focus on enhancing data efficiency through retrieval augmentation. Another line of approach involves compressing the prompt. Although prompt compression methods can be integrated with our approach, either before or after, it would be beneficial to conduct a comparison using the datasets in our study. Specifically, we compare our method with LLMLingua-2 Pan et al. (2024) with a 40% compression rate on the full history (e.g., the actual number of tokens produced can vary due to probabilistic inference). LLMLingua-2 employs a data distillation procedure to extract knowledge from an LLM for prompt compression. The comparative results are presented in Table 6 and Figure 6. Our initial analysis indicates that when comparing the two independent directions of methods separately, our method outperforms in the current setting; this may stem from our method’s integration of knowledge from collaborative users.
@@ A.3 Comparison with Prompt Compression
Table 4: Summary statistics for the RFPN dataset.
A.4 Additional Evaluation Metrics and Qualitative Results
We show results including 1-Wasserstein distance-based alignment score and mean squared error metrics for sentiment intensity prediction in RFPN (Table 5). We see that the results match the trend in the main table.
We additionally show qualitative case studies in Figure 5 to demonstrate the benefits of collaborative strategy.
method wasserstein mse
H-Retrieval 72.77 4.45
H-Recency 72.71 4.44
IntSum 73.15 4.3
History (Full) 73.46 4.3
Persona-DB 74.27 4.13
w/o JOIN 73.13 4.32
w/o IP 72.88 4.49
w/o DP 72.98 4.4
Table 5: Results on additional metrics.
Figure 5: Case Studies.
$\phi_{int}$ (%) $\phi_{p}$ (%)
Method $r_{s}$ $r$ MiF1 MaF1
LLMLingua-2 42.46 42.94 59.96 48.87
Persona-DB 47.67 47.88 62.66 50.59
Table 6: The table displays the RFPN task results when compared to prompt compression.
Figure 6: Comparison with prompt compression on OpinionQA task.
Figure 9: Prompt template used for Baseline Inference.
Figure 11: Prompt template used for Inference for Persona-DB.
Figure 12: Prompt template for retriever.
Figure 13: Prompt template used for Inference for Persona-DB in OpinionQA.